# CDS: Knowledge Component-Driven Data Synthesis Guided by Cognitive Diagnosis Theory

### v2

@@ Abstract
Haokun Zhao1, Jinyi Han2, Jiaqing Liang3 , Yanghua Xiao 1,2 thanks: Corresponding Author Shanghai Key Laboratory of Data Science, School of Computer Science, Fudan University1 Shanghai Institute of Artificial Intelligence for Education, East China Normal University2 Shanghai Key Laboratory of Data Science, School of Data Science, Fudan University3 hkzhao23@m.fudan.edu.cn, haixiahan03@gmail.com, {liangjiaqing,shawyh}@fudan.edu.cn
@@ Appendix A Knowledge Component Details
The following knowledge components (KCs) shown in table 4 are used for various tasks:
@@
Exams Biology: Protein and Nucleic Acid Structure, Sugar and Lipid Types and Functions, Water and Inorganic Salts, Cell Theory, Prokaryotic and Eukaryotic Cells, Cell Membrane Structure, Organelles Structure and Function, Nucleus Structure and Function, Substance Transport Across Cell Membrane, Enzyme Role in Metabolism, ATP Metabolism, Photosynthesis Process, Environmental Impact on Photosynthesis, Cellular Respiration, Cell Growth and Division, Cell Differentiation, Cell Aging and Apoptosis, Cancer Cells and Prevention, Meiotic Division, DNA Structure and Replication, Gene Transcription and Translation, Mendel’s Laws, Sex-linked Inheritance, Gene Mutation, Transgenic Food Safety, Human Genetic Diseases, Evolution Theory, Plant Hormones, Nervous and Hormonal Regulation, Nerve Impulse Transmission, Homeostasis, Immune System Role, Population and Community, Ecosystem Structure and Function, Ecosystem Stability, Biodiversity Conservation, Plant Growth Regulators, Yeast Respiration. Chemistry: Physical vs Chemical Changes, Acids, Bases, Salts and Oxides, Element Symbols, Valency and Formulas, Atomic and Molecular Masses, Law of Mass Conservation, Chemical Reactions, Molar Mass and Volume Calculations, Solubility and Concentration, Colloids and Solutions, Periodic Table Structure, Element Trends, Chemical Bonds, Oxidation-Reduction Reactions, Heat of Reactions, Electrochemistry, Reaction Rate and Activation Energy, Chemical Equilibrium, Electrolytes and Conductivity, pH Calculation, Ionization and Hydrolysis, Organic Compounds, Polymer Chemistry, Laboratory Safety, Gas Production and Separation, Chemical Analysis, Concentration Calculations. Geography: Earth’s Position in Space, Solar Influence, Earth’s Movements and Seasons, Earth’s Layers, Earth Material Cycles, Surface Changes, Atmospheric Heating, Wind and Pressure Systems, Climate and Weather Systems, Water Cycle, Ocean Currents, Geography and Environment, Climate Change, Natural Resources, Natural Disasters, Population Growth, Migration, Urbanization, Agricultural and Industrial Location, Environmental Impact, Geography of Resources, Transportation Systems, Human-Earth Relationships, Sustainability, Green Development, Remote Sensing, Geographic Information Systems, GPS and Navigation, Digital Earth. History: Ancient Chinese Political Systems, Shang and Zhou Dynasties, Qin Centralization, Han to Yuan Political Evolution, Ming and Qing Monarchy, Ancient Chinese Economy, Agricultural Systems, Handicraft and Commerce, Capitalism Emergence, Cultural Evolution, Hundred Schools of Thought, Confucianism, Neo-Confucianism, Chinese Scientific and Technological Achievements, Ancient Greek and Roman Political Systems, Athenian Democracy, Roman Law, Renaissance, Enlightenment, Industrial Revolution, World War Effects, Cold War and Bipolarity, Globalization, WTO and China’s Role, Modern Chinese Politics, Reform and Opening-up, Scientific and Technological Development in China, Modern Chinese Education and Culture. Math: Basic Arithmetic Operations, Decimal and Fraction Operations, Prime and Composite Numbers, Factors and Multiples, GCD and LCM, Algebraic Expressions, Equations, Inequalities, Geometry, Area, Perimeter, Volume, Angles, Coordinates, Mean, Median, Mode, Probability, Permutations, Combinations, Financial Calculations, Unit Conversion, Time and Date Calculations, Speed, Distance, Time, Measurement, Fractions, Decimals, Ratio, Proportion, Bar Graphs, Line Graphs, Probability Theory, Number Sequences, Word Problems, Time Tables, Integer Operations, Linear Equations, Simple Algebra, Pattern Recognition, Mathematical Logic, Shapes and Spatial Understanding, Symmetry, Congruence, Binomial Theorem, Conic Sections. Physics: Motion of Particles, Newton’s Laws, Mechanical Energy, Projectile and Circular Motion, Law of Gravitation, Electric Field, Circuits, Magnetic Field, Electromagnetic Induction, AC, Kinetic Theory of Gases, States of Matter, Thermodynamics, Vibration and Waves, Electromagnetic Waves, Light, Relativity, Reference Frames, Linear Motion, Friction, Hooke’s Law, Scalars and Vectors, Force Composition, Equilibrium, Energy Conservation, Momentum, Collision Theory, Universal Gravitation, Satellite Motion, Electric Potential, Capacitors, Ohm’s Law, Electric Resistance, Magnetic Forces, Lorentz Force, Spectroscopy, Nuclear Physics, Radiation, Energy Loss, Fission and Fusion Reactions, Photoelectric Effect, Einstein’s Equation.
Table 4: Knowledge Components for for various tasks.
@@ D.2 Inference Setup
Data Augmentation temperature=0.5, top_p=0.8, repetition_penalty=1.05, max_tokens=4096, p_rw=0.25, p_fusion=0.25
Data Selection (Stage 1) temperature=0, top_p=1.0, top_k=1, max_tokens=512, repetition_penalty=1.05, $\theta$=8

### v3

@@ Abstract
Haokun Zhao1, Jinyi Han2, Jiaqing Liang3 thanks: Corresponding Author Yanghua Xiao1,2, Xiaojun Meng4, Jiansheng Wei4 1College of Computer Science and Artificial Intelligence, Fudan University 2Shanghai Institute of Artificial Intelligence for Education, East China Normal University 3School of Data Science, Fudan University  4Huawei Noah’s Ark Lab hkzhao23@m.fudan.edu.cn, {liangjiaqing,shawyh}@fudan.edu.cn haixiahan03@gmail.com, {xiaojun.meng,weijiansheng}@huawei.com
@@ Appendix A Knowledge Component Details
The knowledge components (KCs) used for various tasks are shown in Table 4 and Table 5. Table 4 lists the components for Coding and Math, while Table 5 details the components for the various subjects within the Exams task.
@@
Table 4: Knowledge Components for Coding and Math tasks. The components for Exam tasks are in Table 5.
Subject Knowledge Components for Exams
Biology Protein and Nucleic Acid Structure, Sugar and Lipid Types and Functions, Water and Inorganic Salts, Cell Theory, Prokaryotic and Eukaryotic Cells, Cell Membrane Structure, Organelles Structure and Function, Nucleus Structure and Function, Substance Transport Across Cell Membrane, Enzyme Role in Metabolism, ATP Metabolism, Photosynthesis Process, Environmental Impact on Photosynthesis, Cellular Respiration, Cell Growth and Division, Cell Differentiation, Cell Aging and Apoptosis, Cancer Cells and Prevention, Meiotic Division, DNA Structure and Replication, Gene Transcription and Translation, Mendel’s Laws, Sex-linked Inheritance, Gene Mutation, Transgenic Food Safety, Human Genetic Diseases, Evolution Theory, Plant Hormones, Nervous and Hormonal Regulation, Nerve Impulse Transmission, Homeostasis, Immune System Role, Population and Community, Ecosystem Structure and Function, Ecosystem Stability, Biodiversity Conservation, Plant Growth Regulators, Yeast Respiration.
Chemistry Physical vs Chemical Changes, Acids, Bases, Salts and Oxides, Element Symbols, Valency and Formulas, Atomic and Molecular Masses, Law of Mass Conservation, Chemical Reactions, Molar Mass and Volume Calculations, Solubility and Concentration, Colloids and Solutions, Periodic Table Structure, Element Trends, Chemical Bonds, Oxidation-Reduction Reactions, Heat of Reactions, Electrochemistry, Reaction Rate and Activation Energy, Chemical Equilibrium, Electrolytes and Conductivity, pH Calculation, Ionization and Hydrolysis, Organic Compounds, Polymer Chemistry, Laboratory Safety, Gas Production and Separation, Chemical Analysis, Concentration Calculations.
Geography Earth’s Position in Space, Solar Influence, Earth’s Movements and Seasons, Earth’s Layers, Earth Material Cycles, Surface Changes, Atmospheric Heating, Wind and Pressure Systems, Climate and Weather Systems, Water Cycle, Ocean Currents, Geography and Environment, Climate Change, Natural Resources, Natural Disasters, Population Growth, Migration, Urbanization, Agricultural and Industrial Location, Environmental Impact, Geography of Resources, Transportation Systems, Human-Earth Relationships, Sustainability, Green Development, Remote Sensing, Geographic Information Systems, GPS and Navigation, Digital Earth.
History Ancient Chinese Political Systems, Shang and Zhou Dynasties, Qin Centralization, Han to Yuan Political Evolution, Ming and Qing Monarchy, Ancient Chinese Economy, Agricultural Systems, Handicraft and Commerce, Capitalism Emergence, Cultural Evolution, Hundred Schools of Thought, Confucianism, Neo-Confucianism, Chinese Scientific and Technological Achievements, Ancient Greek and Roman Political Systems, Athenian Democracy, Roman Law, Renaissance, Enlightenment, Industrial Revolution, World War Effects, Cold War and Bipolarity, Globalization, WTO and China’s Role, Modern Chinese Politics, Reform and Opening-up, Scientific and Technological Development in China, Modern Chinese Education and Culture.
Math Basic Arithmetic Operations, Decimal and Fraction Operations, Prime and Composite Numbers, Factors and Multiples, GCD and LCM, Algebraic Expressions, Equations, Inequalities, Geometry, Area, Perimeter, Volume, Angles, Coordinates, Mean, Median, Mode, Probability, Permutations, Combinations, Financial Calculations, Unit Conversion, Time and Date Calculations, Speed, Distance, Time, Measurement, Fractions, Decimals, Ratio, Proportion, Bar Graphs, Line Graphs, Probability Theory, Number Sequences, Word Problems, Time Tables, Integer Operations, Linear Equations, Simple Algebra, Pattern Recognition, Mathematical Logic, Shapes and Spatial Understanding, Symmetry, Congruence, Binomial Theorem, Conic Sections.
Physics Motion of Particles, Newton’s Laws, Mechanical Energy, Projectile and Circular Motion, Law of Gravitation, Electric Field, Circuits, Magnetic Field, Electromagnetic Induction, AC, Kinetic Theory of Gases, States of Matter, Thermodynamics, Vibration and Waves, Electromagnetic Waves, Light, Relativity, Reference Frames, Linear Motion, Friction, Hooke’s Law, Scalars and Vectors, Force Composition, Equilibrium, Energy Conservation, Momentum, Collision Theory, Universal Gravitation, Satellite Motion, Electric Potential, Capacitors, Ohm’s Law, Electric Resistance, Magnetic Forces, Lorentz Force, Spectroscopy, Nuclear Physics, Radiation, Energy Loss, Fission and Fusion Reactions, Photoelectric Effect, Einstein’s Equation.
Table 5: Knowledge Components for various Exam subjects. This table details the components for the Exam task mentioned in Table 4.
@@ D.2 Inference Setup
Data Augmentation
temperature=0.5, top_p=0.8, repetition_penalty=1.05, max_tokens=4096,
p_rw=0.25, p_fusion=0.25
Data Selection (Stage 1)
temperature=0, top_p=1.0, top_k=1, max_tokens=512,
repetition_penalty=1.05, $\theta$=8